# Bayesian Ego-graph Inference for Networked Multi-Agent Reinforcement Learning

### v3

@@ 3. Entropy Term: $\log q(Z_{i};\phi_{i})$
$\mathcal{L}_{\mathrm{ELBO}}$ $=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}\right]+\sum_{j\in\mathcal{N}_{i}}\left[\lambda\log\sigma(\phi_{ij})+(1-\lambda)\log(1-\sigma(\phi_{ij}))\right]-\sum_{j\in\mathcal{N}_{i}}H(q(Z_{ij}))$ (32)
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}\right]+\sum_{j\in\mathcal{N}_{i}}\left[\lambda\log\sigma(\phi_{ij})+(1-\lambda)\log(1-\sigma(\phi_{ij}))\right]$
$\qquad\qquad\qquad\qquad\qquad\qquad-\sum_{j\in\mathcal{N}_{i}}\left[-\sigma(\phi_{ij})\log\sigma(\phi_{ij})-(1-\sigma(\phi_{ij}))\log(1-\sigma(\phi_{ij}))\right]$
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\Bigg[-\mathcal{L}_{\theta,\varphi}+\sum_{j\in\mathcal{N}_{i}}\Big(\lambda\log\sigma(\phi_{ij})+(1-\lambda)\log(1-\sigma(\phi_{ij}))$
$\qquad\qquad\qquad\qquad\qquad\qquad+\sigma(\phi_{ij})\log\sigma(\phi_{ij})+(1-\sigma(\phi_{ij}))\log(1-\sigma(\phi_{ij}))\Big)\Bigg]$
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}+\sum_{j\in\mathcal{N}_{i}}\left((\lambda+\sigma(\phi_{ij}))\log\sigma(\phi_{ij})+(2-\lambda-\sigma(\phi_{ij}))\log(1-\sigma(\phi_{ij}))\right)\right],$
@@ Appendix D Algorithms
23:     Update actor: $\theta_{i}\leftarrow\theta_{i}-\eta_{\theta}\nabla\mathcal{L}_{\text{ELBO}}$
24:     Update graph: $\phi_{i}\leftarrow\phi_{i}-\eta_{\phi}\nabla\mathcal{L}_{\text{ELBO}}$
25:   end for
26:   Reset $\mathcal{B}\leftarrow\emptyset$, $k\leftarrow 0$
27:  end if
28:until Stop condition is reached
@@ Appendix E Training Loss Analysis
Figure 10: Qualitative comparison at different simulation times (1000–3599s) on ATSC_Grid. BayesG consistently learns to avoid congestion by adaptively shaping communication, in contrast to fixed or overly dense schemes used by baselines.
@@ Appendix F Additional Visualizations on Grid Environment
Figure 11: Extended case study visualizations for BayesG on ATSC_Grid. Each row shows a different simulation step and training run. (a) Left: Latent interaction matrix (ego-graph probabilities). (a) Right: Physical network view with edge thickness indicating communication strength. (b): Corresponding vehicle congestion snapshot from SUMO.

### v4

@@ 3. Entropy Term: $\log q(Z_{i};\phi_{i})$
$\mathcal{L}_{\mathrm{ELBO}}$ $=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}\right]+\sum_{j\in\mathcal{N}_{i}}\left[\sigma(\phi_{ij})\log\lambda+(1-\sigma(\phi_{ij}))\log(1-\lambda)\right]+\sum_{j\in\mathcal{N}_{i}}H(q(Z_{ij}))$ (36)
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}\right]+\sum_{j\in\mathcal{N}_{i}}\left[\sigma(\phi_{ij})\log\lambda+(1-\sigma(\phi_{ij}))\log(1-\lambda)\right]$
$\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{j\in\mathcal{N}_{i}}\left[-\sigma(\phi_{ij})\log\sigma(\phi_{ij})-(1-\sigma(\phi_{ij}))\log(1-\sigma(\phi_{ij}))\right]$
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\Bigg[-\mathcal{L}_{\theta,\varphi}+\sum_{j\in\mathcal{N}_{i}}\Big(\sigma(\phi_{ij})\log\lambda+(1-\sigma(\phi_{ij}))\log(1-\lambda)$
$\qquad\qquad\qquad\qquad\qquad\qquad-\sigma(\phi_{ij})\log\sigma(\phi_{ij})-(1-\sigma(\phi_{ij}))\log(1-\sigma(\phi_{ij}))\Big)\Bigg]$
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}+\sum_{j\in\mathcal{N}_{i}}\left(\sigma(\phi_{ij})\log\frac{\lambda}{\sigma(\phi_{ij})}+(1-\sigma(\phi_{ij}))\log\frac{1-\lambda}{1-\sigma(\phi_{ij})}\right)\right]$
$=\mathbb{E}_{q(Z_{i};\phi_{i})}\left[-\mathcal{L}_{\theta,\varphi}\right]-\sum_{j\in\mathcal{N}_{i}}\mathrm{KL}\left(q(Z_{ij};\phi_{ij})\|p(Z_{ij})\right),$
@@ Appendix D Algorithms
23:     Update actor & graph: $\theta_{i},\phi_{i}\leftarrow\theta_{i},\phi_{i}-\eta_{\theta}\nabla_{\theta_{i},\phi_{i}}(-\mathcal{L}_{\text{ELBO}})$
24:   end for
25:   Reset $\mathcal{B}\leftarrow\emptyset$, $k\leftarrow 0$
26:  end if
27:until Stop condition is reached
@@ Appendix E Training Loss Analysis
Figure 10: Qualitative comparison at different simulation times (1000–3599s) on ATSC_Grid. Road thickness and color represent vehicle density (thicker and redder indicates more congestion). BayesG consistently learns to avoid congestion by adaptively shaping communication, in contrast to fixed or overly dense schemes used by baselines.
@@ Appendix F Additional Visualizations on Grid Environment
Figure 11: Extended case study visualizations for BayesG on ATSC_Grid. Each row shows a different simulation step and training run. (a) Left: Latent interaction matrix (ego-graph probabilities). (a) Right: Physical network view with edge thickness indicating communication strength.